OpenStudy (anonymous):

Medal and fan! The sum of two numbers is 48. If one third of one number is 5 greater than one sixth of another number, which of the following is the smaller number? 6 26 22 42

OpenStudy (anonymous):

n1 = 48 - n2 1/3 n1 = 1/6n2 + 5 So 1/3 (48-n2) = 1/6n2 + 5 16 - 1/3n2 = 1/6n2 + 5 11 = 1/6n2 + 1/3n2 11 = 1/2n2 n2 = 22

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Np